# An upper bound of the silhouette validation metric for clustering

### v2

@@ 5 Experiments, results & discussion
Table 1: Real-world datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
@@
Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound.
Dataset $K$ ASW $\texttt{UB}(\Delta)$ Worst-case rel. error
400-64-5-6 5 $0.249$ $0.376$ $0.339$
400-64-2-2 2 $0.673$ $0.673$ $0.000$
400-128-7-3 7 $0.522$ $0.566$ $0.078$
1000-300-5-2 5 $0.668$ $0.678$ $0.015$
10000-32-20-2 20 $0.626$ $0.774$ $0.192$
10000-1024-20-4 20 $0.417$ $0.454$ $0.081$
We evaluate the ASW upper bound by applying it to a variety of datasets, both synthetic and real-world. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
An essential part of the evaluation is to establish a nearly tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we use PAMSIL clustering to obtain an empirical $\mathrm{ASW}$. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies. PAMSIL is a suitable algorithm since its objective is to maximize ASW.
Given an empirical ASW value, say $\widehat{S}$, the performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes a clustering that achieves the global maximum. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical ASW is guaranteed to be within $\epsilon$ of the true maximum.
Synthetic datasets
We follow [gaussian] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [scikit]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion. Finally, for these datasets, we use Euclidean distance.
Real-world datasets
We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [ucidata], commonly used in clustering and classification research. Table 1 summarizes their key characteristics along with the distance metrics employed. Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate. Each dataset provides a known number of classes, which we use to determine the number of clusters $K$. When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures. In that way, we ensure that our experiments are conducted in natural clustering contexts.
Table 3: Experimental results on real data. Comparison of clustering performance and the proposed upper bound.
Performance of PAMSIL clustering
Comparison with
global upper bound
Comparison with
adjusted upper bound
Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI ASW $\texttt{UB}(\Delta)$ Worst-case rel. error $\texttt{UB}_{\kappa}(\Delta)$ Worst-case rel. error
@@
Figure 5: Silhouette plots for the eight real-world datasets. PAMSIL clustering with $K$ equal to the ground truth number of classes; ASW compared against the global upper $\texttt{UB}(\Delta)$ and the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
@@
Results
In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering. For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score. For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL ASW is within 8 $\%$ of the global maximum, that is, essentially optimal. For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical ASW of $0.249$. In this case, it remains unclear whether further substantial improvements of the ASW are possible. However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
Table 3 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds. In these cases, the global upper bounds provide information that is less actionable. It is unclear whether the reason for this is that the attained ASW values are far from $\mathrm{ASW}(C^{*},\Delta)$, or that the upper bounds are somewhat loose, or both. However, the adjusted upper bounds $\texttt{UB}_{\kappa}(\Delta)$ are considerably tighter for most datasets. Here, $\kappa$ is set to the minimum cluster size in the PAMSIL solution. For datasets Ceramic, Customers, Rna, Wdbc and Wine, $\texttt{UB}_{\kappa}(\Delta)$ proves that the ASW-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums.
@@ Results
Fig. 5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds. For the adjusted upper bounds, $\kappa$ is set to the smallest cluster size in the generated solution. This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
@@
In Fig. 6 we show, for each real-world dataset and for each $K\in\{2,\ldots,15\}$, the attained ASW value compared to the global ASW upper bound. Additionally, at each $K$, we compare the ASW to the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$ with $\kappa$ set to the minimum cluster size in the corresponding solution. Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2.
Discussion
We summarize with some reflections on the practical relevance of the proposed upper bound.
@@ Discussion
Replacing the standard $\mathrm{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\mathrm{ASW}$ values. However, when the empirical $\mathrm{ASW}$ lies far below the upper bound, it remains unclear whether further improvement is actually possible. We suspect that this gap often persists for the global upper bound, even when the empirical $\mathrm{ASW}$ is close to $\mathrm{ASW}(C^{*},\Delta)$. In many real-world applications, however, analysts impose minimum cluster-size constraints to avoid overly small or degenerate clusters. By incorporating such constraints through $\kappa$, the refined bound $\texttt{UB}_{\kappa}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice. As demonstrated in Table 3, $\texttt{UB}_{\kappa}(\Delta)$ ensures that the PAMSIL solutions are within 30 $\%$ of the optimum for five of the examined datasets, indicating strong empirical performance under realistic conditions. In such cases, the bound can be used to conclude that further exploration would, at best, yield only marginally better results.
6 Conclusion
@@
In this paper, we proposed a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a complementary tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value. It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often nearly tight and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should not be viewed as a precise target but as an informative ceiling that frames observed ASW scores in a sharper context.
@@
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices. An important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its tightness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters. Such insights would clarify in which contexts the bound provides the greatest practical value.
@@
Hugo Sträng: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: visualization, writing, editing, validation.

### v4

@@ 5.2 UCI datasets
Table 3: UCI datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
@@
Table 4: Experimental results on UCI datasets. Comparison of clustering performance and the proposed upper bound.
Performance of PAMSIL clustering Comparison with global upper bound Comparison with constrained upper bound
Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$
@@
Table 4 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds. In these cases, the gap between the empirical $\operatorname{ASW}$ and the global upper bound is generally quite large. It is unclear whether the reason for this is that the attained $\operatorname{ASW}$ values are far from $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$, or that the upper bounds are rather loose, or both. However, the constrained upper bounds $\texttt{UB}_{m}(\Delta)$ are considerably tighter for most datasets. Here, $m$ is set to the minimum cluster size in the PAMSIL solution. For datasets Ceramic, Customers, Rna, Wdbc, and Wine, $\texttt{UB}_{m}(\Delta)$ proves that the $\operatorname{ASW}$-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums (within the solution spaces constrained by $m$).
Figure 5: Silhouette plots for the UCI datasets. PAMSIL clustering with $K$ equal to the ground truth number of classes; $\operatorname{ASW}$ compared against the global upper $\texttt{UB}(\Delta)$ and the constrained upper bound $\texttt{UB}_{m}(\Delta)$, with $m$ being the smallest cluster size. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding constrained upper bounds, while the transparent areas indicate the corresponding global upper bounds.
@@
As for the global version, the lowest upper bound we recorded was for the Rna dataset, for which PAMSIL reached an ASW of $0.225$ and the upper bound was $0.419$. In contrast, the highest upper bound we recorded was for the Customers dataset, for which PAMSIL reached an ASW of $0.4$ and the upper bound was $0.951$. The smallest and largest worst-case relative errors were for the Ceramic dataset ($0.423$) and the Heart Statlog dataset ($0.731$), respectively.
We observe that the gaps between $\texttt{UB}(\Delta)$ and $\texttt{UB}_{m}(\Delta)$ are rather large for most datasets. For instance, the Heart Statlog dataset showed $\texttt{UB}(\Delta)=0.6$ and $\texttt{UB}_{83}(\Delta)=0.28$ — a 53 $\%$ decrease. Earlier in the text, when discussing Figure 3, we speculated that, in many real-world datasets, the $k$-quotient will frequently reach its minimum at $k=2$ regardless of the cluster structure in $\mathcal{C}^{*}$. If $q(\Delta,m)\gg q(\Delta,2)$ this naturally leads to large gaps. On the one hand, the constrained upper bound is an important feature of the proposed method that can lead to significantly sharper insights than those provided by the global upper bound. On the other hand, implementing an unjustified constraint by mistake can cause incorrect conclusions. Therefore, we recommend careful consideration when choosing the parameter $m$.
@@
Figure  5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding constrained upper bounds, while the transparent areas indicate the corresponding global upper bounds. For the constrained upper bounds, $m$ is set to the smallest cluster size in the generated solution. This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
@@
In Figure  6 we show, for each dataset and for each $K\in\{2,\ldots,15\}$, the attained $\operatorname{ASW}$ value compared to the global $\operatorname{ASW}$ upper bound. Additionally, at each $K$, we compare the $\operatorname{ASW}$ to the constrained upper bound $\texttt{UB}_{m}(\Delta)$ with $m$ set to the minimum cluster size in the corresponding solution. Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2. This scan over different $K$ reveals that PAMSIL reaches a higher or equally high $\operatorname{ASW}$ at a value of $K$ that is different than the ground truth number of classes for all datasets but Ceramic and Wdbc.
5.3 The ALOI dataset
The purpose of this experiment is to study how the proposed upper bound behaves on larger datasets with a large number of clusters. In particular, we investigate whether the bound appears informative when the number of classes is large.
We base our experiments on data derived from the Amsterdam Library of Object Images (ALOI) [21], a large-scale image collection containing objects photographed under varying viewing angles and illumination conditions. Each image is represented using HSV/HSB color histograms with different binning resolutions. In this representation, each pixel is transformed into hue–saturation–brightness space and assigned to its nearest histogram bin. We consider two such representations, namely $5\!\times\!5\!\times\!5$ and $14\!\times\!6\!\times\!6$ binnings, resulting in feature vectors of up to several hundred dimensions. In this setting, we use the Manhattan distance metric.
The datasets have 1000 object classes with at least 100 instances available for each class. To obtain a total number of samples that is manageable, we randomly select $40$ instances per class, yielding datasets with 40,00040{,}000 samples in total. In this regime, PAMSIL becomes computationally prohibitive, and we therefore rely on FastMSC and DynMSC, which are significantly more scalable while still directly optimizing silhouette-based objectives. For these experiments, we report both clustering runtimes and the runtimes required to compute the global and constrained upper bounds. Table 5 summarizes the results.
With $K=1000$, the gap between the global upper bound and the empirical $\operatorname{ASW}$ is rather substantial for both datasets. However, applying DynMSC ($2\leq K\leq 50$) reveals that a much higher $\operatorname{ASW}$ is achievable at $K=2$.
For the $14\!\times\!6\!\times\!6$ dataset with $m=2$ and $m=3$, the constrained upper bounds show slightly higher runtimes than the global upper bounds, which we attribute to runtime variability in the experimental environment.
We also conduct some small-scale experiments with 2 and 5 classes from the same underlying data. For these datasets, we sample 100 instances from each class to obtain datasets of sizes 200 and 500, respectively. For instance, $5\!\times\!5\!\times\!5-2$ and $5\!\times\!5\!\times\!5-5$ are datasets derived from the same image representations, but with different number of classes (2 and 5). Since these datasets are so small, we can apply PAMSIL and compare it with FastMSC and DynMSC. Table 6 summarizes the results.
It is difficult to draw strong conclusions from Table 5. The large-scale setting introduces substantial uncertainty regarding the most appropriate number of clusters, which complicates a precise assessment of the bounds tightness.
In contrast, Table 6 provides a more controlled comparison. The results suggest that the gap between the maximum $\operatorname{ASW}$ and the proposed upper bound increases as the number of clusters in the desirable solution grows, even when moving from just $2$ to $5$ clusters. This observation highlights an important limitation of the method: the bound appears to be most informative when the underlying clustering structure involves a relatively small number of clusters. Consequently, the proposed upper bound should be interpreted with great caution in settings where the true number of clusters is large, and its practical usefulness is expected to be greatest when a small number of clusters is assumed.
Table 5: Experimental results on ALOI datasets with 1000 classes. Comparison of clustering performance and the proposed upper bound. DynMSC scans $2\leq K\leq 50$ and FastMSC scans $K=100,200,\ldots,900$. Runtime (RT) is expressed in 1000 seconds.
Performance of FastMSC and DynMSC clustering Comparison with global upper bound Comparison with constrained upper bound
Dataset Algorithm $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ RT $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ RT $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$ RT
5x5x5 FastMSC 1000 2 $0.267$ $0.557$ $0.133$ $1.1$ $0.914$ $0.854$ $0.3$ $0.914$ $0.854$ $0.3$
FastMSC 100 7 $0.006$ $0.368$ $0.181$ $9.7$ $0.914$ $0.802$ $0.3$ $0.866$ $0.791$ $0.3$
DynMSC 2 976 $0.000$ $0.009$ $0.502$ $1.6$ $0.914$ $0.451$ $0.3$ $0.618$ $0.188$ $0.3$
14x6x6 FastMSC 1000 2 $0.292$ $0.589$ $0.142$ $1.1$ $0.901$ $0.842$ $0.2$ $0.901$ $0.842$ $0.3$
FastMSC 100 3 $0.005$ $0.352$ $0.161$ $9.9$ $0.901$ $0.822$ $0.2$ $0.883$ $0.818$ $0.3$
DynMSC 2 38 $0.000$ $0.001$ $0.470$ $1.7$ $0.901$ $0.478$ $0.2$ $0.752$ $0.375$ $0.2$
Table 6: Experimental results on ALOI datasets with a small number of classes. Comparison of clustering performance and the proposed upper bound. DynMSC scans $2\leq K\leq 10$.
Performance of PAMSIL clustering Comparison with global upper bound Comparison with constrained upper bound
Dataset Algorithm $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$
5x5x5-2 PAMSIL 2 94 $0.883$ $0.834$ $0.669$ $0.899$ $0.256$ $0.715$ $0.064$
FastMSC 2 94 $0.883$ $0.834$ $0.669$ $0.899$ $0.256$ $0.715$ $0.064$
DynMSC 8 1 $0.876$ $0.820$ $0.647$ $0.899$ $0.281$ $0.899$ $0.281$
5x5x5-5 PAMSIL 5 24 $0.675$ $0.751$ $0.478$ $0.899$ $0.468$ $0.762$ $0.373$
FastMSC 5 19 $0.654$ $0.723$ $0.473$ $0.899$ $0.473$ $0.778$ $0.392$
DynMSC 2 28 $0.001$ $0.009$ $0.416$ $0.899$ $0.537$ $0.752$ $0.446$
14x6x6-2 PAMSIL 2 94 $0.883$ $0.834$ $0.640$ $0.879$ $0.271$ $0.695$ $0.078$
FastMSC 2 92 $0.846$ $0.795$ $0.639$ $0.879$ $0.273$ $0.702$ $0.090$
DynMSC 9 1 $0.856$ $0.776$ $0.644$ $0.879$ $0.267$ $0.879$ $0.267$
14x6x6-5 PAMSIL 5 24 $0.670$ $0.739$ $0.473$ $0.888$ $0.467$ $0.752$ $0.371$
FastMSC 5 23 $0.676$ $0.754$ $0.471$ $0.888$ $0.470$ $0.755$ $0.376$
DynMSC 6 20 $0.678$ $0.746$ $0.472$ $0.888$ $0.468$ $0.764$ $0.382$
5.4 Discussion
In Table 2, we saw that $\texttt{UB}(\Delta)$ was able to confirm that PAMSIL reached a globally optimal $\operatorname{ASW}$ for the synthetic dataset 400-64-2-2. This is interesting from a mathematical perspective, but it is not a realistic use case. For any real-world dataset, we expect the likelihood of the bound being sharp to be rather low.
@@ 5.4 Discussion
Replacing the standard $\operatorname{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\operatorname{ASW}$ values. However, when the empirical $\operatorname{ASW}$ lies far below the upper bound, it remains unclear to what extent further improvement is actually possible. We suspect this gap to often be substantial for the global upper bound, even when the empirical $\operatorname{ASW}$ is close to $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$. In many real-world applications, however, analysts impose minimum cluster size constraints to avoid overly small or degenerate clusters [24, Höppner2008]. By incorporating such constraints through $m$, the refined bound $\texttt{UB}_{m}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice. As demonstrated in Table 4, $\texttt{UB}_{m}(\Delta)$ ensures that the PAMSIL solution is within 30 $\%$ of the optimum (in the constrained solution space) for five of the examined datasets.
The experimental results indicate that the usefulness of the upper bound varies between different contexts; understanding when it is most beneficial in practice is important. For the sake of argument, we can assume that the bound is useful whenever it is far from $1.0$ (say, less than $0.5$). Following this logic, one can expect the bound to be most informative when the dimensionality of the data is rather high. The curse of dimensionality tells us that distances become more similar when dimensionality increases, which should result in the upper bound getting closer to zero. Indeed, for the real-world cases, the smallest upper bound we observed was for the Rna dataset with Euclidean distances and 20531 features. In contrast, the largest upper bound we observed was for the Customers dataset (Cosine distance), with only 7 features. In addition to the fact that the data should be high-dimensional, we speculate that the upper bound is more useful when the number of clusters in the optimal solution is relatively low.
6 Summary
6.1 Limitations
Challenges with the ASW
A natural limitation of the proposed method stems from the limitations of the silhouette metric itself. Many clustering quality indices have been proposed over time, and the ASW is not inherently superior to these. Its usefulness depends on the structure of the data. For instance, the ASW can be misleading when the underlying clusters are of different diameters or strongly anisotropic, in which case a desirable partition can even result in negative silhouette widths. Also, singletons are usually given a silhouette width of 0 by default, even when they are well separated. In conclusion, the proposed upper bound should only be considered in situations where the silhouette is meaningful.
Non-sharpness
The proposed upper bound is not guaranteed to be close to the maximum ASW. Therefore, the upper bound should not be viewed as a precise target, but rather an informative ceiling. Generally, we expect the bound to be rather loose when the number of underlying clusters is large.
Challenges with evaluation
A fundamental limitation inherent to the problem is that the true global maximum of the ASW is unknown for any non-trivial dataset. Consequently, the quality of our upper bound cannot be assessed exactly, but only indirectly through empirical evidence. This makes it difficult to draw precise conclusions.
Poor scalability
Since the proposed algorithm assumes the full dissimilarity matrix as input, it suffers from the same scalability issue as that for computing the ASW, namely that $\mathcal{O}(n^{2})$ memory is required. In principle, one could design an implementation that instead computes and sorts distances per data point, reducing the memory requirement to $\mathcal{O}(n)$. However, this does not change the overall time complexity, which remains $\mathcal{O}(n^{2}\log n)$. Therefore, the proposed method does not scale to very large datasets.
Hardware constraints
Due to hardware constraints, large datasets (e.g., $n>50000$) were not included in our empirical evaluation. This does not affect the theoretical correctness of the method, but limits our empirical assessment at large scales.
6.2 Conclusion
@@
In this paper, we proposed a dataset-dependent upper bound on the $\operatorname{ASW}$ as a complementary tool for analyzing clustering quality. While $\operatorname{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more relevant reference value. It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
As for the practical usefulness of the proposed upper bound, our findings do not indicate that meaningful insights can be expected in the general case. However, it is evident from the results that there are contexts in which the upper bound succeeds in providing useful information. For example, for the Rna dataset, we observed an upper bound of $0.42$, which placed the empirical $\operatorname{ASW}$ of $0.23$ in a completely different lighting. Pinpointing these cases where the upper bound is the most beneficial would be of great value.
@@
From a broader perspective, our main contribution has been to establish a proof of concept: that efficiently computable, data-dependent upper bounds on $\operatorname{ASW}$ can enrich cluster analysis. We believe that this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices, for example, to the medoid-based silhouette. An important open theoretical question, which we do not address here, is to characterize how loose the proposed upper bound is in a given case, for example, by deriving a lower bound on the gap between the true maximum $\operatorname{ASW}$ and $\texttt{UB}(\Delta)$. As mentioned above, another important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its usefulness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters. Such insights would clarify in which contexts the bound provides the greatest practical value.
@@
Hugo Sträng: Conceptualization, Methodology, Investigation, Data Curation, Writing - Original Draft, Writing - Review &\& Editing, Visualization; Tai Dinh: Validation, Writing - Review &\& Editing, Visualization.